# Autonomous Robotic Bone Micro-Milling System with Automatic Calibration and 3D Surface Fitting

### v2

@@ III-B1 Trajectory Center Localization
III-B2 3D-fitted Initial Trajectory Generation
Considering the expected trajectory is circular when projected onto the $x\text{--}y$ plane, by combining the obtained trajectory center $\boldsymbol{p}_{c}^{\{\text{Mi}\}}$ and the radius of the milling trajectory $r$, the $x$ and $y$ coordinates of $n$ sampling points along the trajectory $\boldsymbol{p}_{fit,i}^{\{\text{Mi}\}}\triangleq\begin{bmatrix}x_{fit,i}&y_{fit,i}&z_{fit,i}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{3},i=1,\cdots,n\in\mathbb{N}$, with respect to {Mi}, can be calculated by:
@@
where $\theta_{i}=2{\pi}i/n,i=1,\cdots,n\in\mathbb{N}$. Then $z_{fit,i}$ can be obtained by combining $x_{fit,i},\ y_{fit,i}$ and the reconstructed point cloud of the target surface.
In order to avoid contact with the surface even before milling, an offset should be set for all sampling points along their normal directions, considering the radius of the drill tip $r_{d}$ and the potential error of Automatic Calibration. The normal vectors of all sampling points $\boldsymbol{n}_{i}\in\mathbb{R}^{3},i=1,\cdots,n\in\mathbb{N}$ can also be obtained by the reconstructed point cloud of the target surface. As a result, by substituting the result of (5), the 3D-fitted initial trajectory $\boldsymbol{p}_{ini}^{\{\text{Mi}\}}$ with respect to {Mi} can be obtained, whose each element can be calculated by
@@
where $\Delta d$ is a manually set offset for safety concerns, considering errors in Automatic Calibration and 3D Surface Fitting.
@@
The result of (7) is output to the Trajectory Planner (Section IV-B) for the initialization of the milling trajectory.
IV End-to-end Automation Workflow
With the integration of the proposed pre-measurement pipeline (Section III), we improve the autonomous robotic bone micro-milling system and achieve an end-to-end automation workflow (see Fig. 2). In this section, we briefly introduce the other modules of the workflow, namely Completion Level Recognition (Section IV-A) and Trajectory Planner (Section IV-B). These modules are modified from the existing system in the authors’ prior research [8] and play an essential role in providing real-time feedback after the pre-measurement phase is completed and milling begins.
IV-A Completion Level Recognition
The Completion Level Recognition aims to monitor the milling completion level along the trajectory in real-time using a CNN based on the input RGB images. The CNN follows the same DSSD-inspired [17] architecture as the previous system [8], but is retrained on a new dataset to adapt to the mouse skull milling experiment. The CNN output a bounding box for milling area detection and a heatmap image for pixel-wise completion level prediction. The network was trained on an augmented dataset derived from 587 manually collected and annotated images (a total of 18,784 training images) from videos of teleoperated robotic cranial window creation experiments conducted on 6 euthanized mice, achieving 77.6 mAP for the detection task and 24.32% MAPE for the completion prediction task, with a processing speed of 72 Hz. More network details are available in the supplementary document.
Through post-processing and sampling of the output image, an $n$-dimensional vector $\boldsymbol{c}$ can be obtained, in which element denotes the penetration percentage at the $i$-th sampling point along the real-time-updated trajectory, defined by
@@ IV-A Completion Level Recognition
where $i=1,2,...,n$. Note that $c_{i}=0$ means untouched and $c_{i}=1$ means completely penetrated. $\boldsymbol{c}$ is output to the Trajectory Planner for real-time trajectory updates.
IV-B Trajectory Planner
@@
The Trajectory Planner is responsible for initializing, updating and smoothing the milling trajectory. The updating process remains consistent with the previous system, employing a multimodal velocity damper to control the descent speed of the drill bit at the $i$-th sampling point to proportionally reduce based on its completion level $c_{i}$, which is real-time output from the Completion Level Recognition and defined by (8).
@@
For the initialization process, in the previous system, due to the absence of direct surface measurements, the $z$-coordinates of all sampling points are initialized to 0 (see Fig. 5-(a)). In this study, with the integration of MSCS, a 3D-fitted initial trajectory $\boldsymbol{p}_{ini,i}$ is available, which is output from the 3D Surface Fitting (Section III-B) and defined by (1). The trajectory features non-zero $z$-axis coordinate at each sampling point to adapt to the uneven surface for (see Fig. 5-(b)), enabling the drill to contact the target surface rapidly and spatially consistently.
@@
The smoothing process also remains the same as in the previous system, where a constrained cubic spline interpolation method [18] is used to generate a continuous trajectory based on the $n$ discrete sampling points, to match the high-precision encoders used in the robotic system [9]. Additionally, we hypothesize that if all $n$ sampling points are completely penetrated, the entire smoothed trajectory will be sufficiently milled. Therefore, when $c_{i}=1\ \ \forall i=1,\cdots,n\in\mathbb{N}$, a stop signal will be triggered and sent to the controller to automatically halt the robot’s movement and the drill’s rotation, marking the completion of the milling process. More details of the Trajectory Planner are available in the supplementary document.
V Experiments
@@
In this section, experiments are first conducted to validate the effectiveness of the Automatic Calibration and the 3D Surface Fitting, which together constitute the proposed pre-measurement pipeline. Subsequently, further experiments are conducted to evaluate the performance of the entire autonomous robotic bone micro-milling system.
V-A System configuration
The software implementation of the experiments was carried out on an Ubuntu 20.04. The robotic arm was controlled following the approach in [9]. ROS (Robot Operating System) Noetic Ninjemys was utilized for interprocess communication, while simulations were conducted using CoppeliaSim (Coppelia Robotics, Switzerland). Communication with the robot was managed via the SmartArmStack [^2] https://github.com/SmartArmStack. Dual quaternion algebra and robot kinematics were implemented using DQ Robotics [19] in Python 3.
V-B Parameter selection
@@
To implement the proposed method for validations and experiments, the parameters mentioned in Section III and Section IV are selected based on camera calibration, actual measurements, safety considerations, and ease of calculation: the number of trajectory sampling points $n=32$, the intrinsic parameters of cameras $c_{x}=480\ \mathrm{pixel}$, $c_{y}=270\ \mathrm{pixel}$, $P_{\rho x}=0.03947\ \mathrm{pixel/mm}$ , $P_{\rho y}=0.03894\ \mathrm{pixel/mm}$, $d_{e}=0$, and $h_{\rho}=6.4\ \mathrm{pixel/mm}$, the ratio of the distance between the trajectory center and bregma to the distance between bregma and lambda $l=1/3$, the radius of the milling trajectory $r=2\ \mathrm{mm}$, the radius of the drill tip $r_{d}=0.5\ \mathrm{mm}$, the manually set offset $\Delta d=0.5\ \mathrm{mm}$, the number of calibration setpoints $k=5$ and their coordinates $\boldsymbol{p}_{1}^{\{\text{R}\}}=\begin{bmatrix}0&-20&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{2}^{\{\text{R}\}}=\begin{bmatrix}10&-20&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{3}^{\{\text{R}\}}=\begin{bmatrix}5&-10&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{4}^{\{\text{R}\}}=\begin{bmatrix}10&0&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{5}^{\{\text{R}\}}=\begin{bmatrix}0&0&0\end{bmatrix}^{\text{T}}\ \mathrm{mm}$. Additionally, the resolution of the images captured by the MSCS is set to $540\times 960$ to meet the requirement for Automatic Calibration, 3D Surface Fitting, and Completion Level Recognition.
V-C Experiment 1: Validation of Automatic Calibration
@@
As described in Section III-A, the accuracy of Automatic Calibration is affected by the accuracy of 2D Keypoint Detection, 3D Reconstruction, and Landmark Calibration. Considering the 3D Reconstruction depends on the MSCS, whose accuracy has been evaluated in the prior research by the present authors [11], in Experiment 1, we focus on the evaluation of 2D Keypoint Detection and Landmark Calibration.
V-C1 Training of 2D Keypoint Detection model
@@
To prepare learning data, a total of 3,294 images satisfying the requirement that the drill tip is within the field of view were captured from videos of teleoperated robotic cranial window creation experiments conducted on 10 euthanized mice. The mice were sacrificed for another research and reused in this study, so no additional ethical approval was required. For each image, a confidence map is generated by applying a 2D Gaussian distribution to the manually annotated drill tip 2D position, which represents the probability of the drill tip’s existence and serves as the ground truth (an example is shown in Fig. 6-(a)). Data augmentation techniques, including blur and noise augmentation, color adjustment, geometric transformation, and resolution alteration, resulted in a total of 16,394 images and corresponding ground truths, with 14,754 images used for training, 820 for validation, and 820 for testing.
@@
The models were trained in a total of 30 epochs using the Adam optimizer, with the training loss computed based on the Mean Square Error (MSE) criterion. The training rate was initialized as $1\times 10^{-3}$, followed by a reduction to $1\times 10^{-4}$ in epochs 11 to 20, and further decreased to $1\times 10^{-5}$ for epochs 21 to 30. Moreover, a batch size of 4 was employed during the training phase. All models were implemented using PyTorch 1.8.0 with CUDA 11.1, based on Python 3.8, and executed on an Ubuntu 20.04 system equipped with an NVIDIA Quadro P6000 graphics card.
V-C2 Joint evaluation metrics
@@
Firstly, we send the robot $m=20$ validation setpoints $\boldsymbol{p}_{val,i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,m$ for the drill tip to reach these points. MSCS captures $m$ images each from left and right camera, outputting $m$ drill tip 2D positions each $\boldsymbol{p}_{l,i},\ \boldsymbol{p}_{r,i}\in\mathbb{R}^{2},i=1,\cdots,m$ via 2D Keypoint Detection (see Section III-A1). Then we manually annotate the left and right images, obtaining $m$ ground truths each $\boldsymbol{p}_{l,i}^{(gt)},\ \boldsymbol{p}_{r,i}^{(gt)}\in\mathbb{R}^{2},i=1,\cdots,m$. Through 3D Reconstruction (see Section III-A2), $m$ points of 3D reconstructed drill tip with respect to {Mi} can be obtained. Then, by transforming these points back to {R} using the transformation calculated from Landmark Calibration (see Section III-A3), we can obtain $\boldsymbol{p}_{trans,i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,m$. The accuracies of 2D Keypoint Detection and Landmark Calibration are respectively evaluated by the Root Mean Square Error (RMSE) between $\boldsymbol{p}_{l,i},\ \boldsymbol{p}_{r,i}$ and $\boldsymbol{p}_{l,i}^{(gt)},\ \boldsymbol{p}_{r,i}^{(gt)}$, and between $\boldsymbol{p}_{trans,i}^{\{\text{R}\}}$ and $\boldsymbol{p}_{val,i}^{\{\text{R}\}}$.
Given the image resolution of $540\times 960$, the error in 2D Keypoint Detection is required to not exceed 10 pixels (0.04 mm). Considering the radius of the drill tip ($r_{d}=0.5\ \mathrm{mm}$) and the manually set offset ($\Delta d=0.5\ \mathrm{mm}$) in (6), the errors of Landmark Calibration is required to not exceed 1 mm.
V-C3 Result and discussion
@@
As shown in Fig. 7, we randomly generated 20 coordinates as validation setpoints within the reconstruction space of MSCS, and observed an RMSE of 3.18 pixels in 2D Keypoint Detection and 0.77 mm in Landmark Calibration, which meet our accuracy requirement, indicating a high overall detection and calibration precision. Specifically, as shown in Fig. 7-(a), all 20 validation points exhibited 2D Keypoint Detection errors within the threshold, indicating that the detection model demonstrated good accuracy and robustness. However, as shown in Fig. 7-(b), 4 points exceeded the error threshold in Landmark Calibration, which is likely due to the instability of the MSCS reconstruction accuracy in certain spatial regions, such as areas with sparse textures, suboptimal viewpoints, or significant depth variations. Although these localized errors are still acceptable in Experiment 1, they may lead to milling failures in the subsequent Experiment 3. To address this issue, future work will focus on improving 3D reconstruction accuracy by employing higher-resolution cameras and more robust models, thereby enhancing spatial localization performance.
V-D Experiment 2: Validation of 3D Surface Fitting
As described in Section III-B, the accuracy of 3D Surface Fitting is affected by the accuracy of Trajectory Center Localization and 3D-fitted Initial Trajectory Generation. As the performance of 3D-fitted Initial Trajectory Generation is based on the reconstruction of the target surface, which has been evaluated in [11], and the transformation obtained via Landmark Calibration, which has been evaluated in Section V-C, in Experiment 2, we focus on the evaluation of Trajectory Center Localization.
V-D1 Training of CNN model
To prepare learning data, a dataset consisting of 523 images that satisfy the requirement that the mouse skulls in the images remain unmilled was captured from videos of the same 10 euthanized mice mentioned in Section V-C, as well as 82 images from a publicly available dataset created by [16]. The creation of ground truths and data augmentation are the same as Section V-C, resulting in a total of 8,774 images, with 7,019 for training, 877 for validation, and 878 for testing (an example is shown in Fig. 6-(b)). The training condition and parameters are also as same as Section V-C.
V-D2 Metrics
Considering that the trajectory center is calculated by linearly weighting the detected bregma and lambda, with the detection errors of these two points considered independent, the error of Trajectory Center Localization can also be calculated by linearly weighting the detection RMSE of bregma and lambda. Same as Section V-C, we require the error not to exceed 10 pixels.
V-D3 Result and discussion

### v3

@@ II-B1 Trajectory Center Localization
II-B2 3D-fitted Initial Trajectory Generation
Considering that the expected trajectory is circular when projected onto the $x\text{--}y$ plane, by combining the obtained trajectory center $\boldsymbol{p}_{c}^{\{\text{Mi}\}}$ and the milling trajectory radius $r$, a surface-fitted trajectory $\boldsymbol{p}_{fit}^{\{\text{Mi}\}}$ with respect to {Mi} can be sampled on the reconstructed point cloud. The $x$ and $y$ coordinates of its $i$-th sampling point $\boldsymbol{p}_{fit,i}^{\{\text{Mi}\}}\triangleq\begin{bmatrix}x_{fit,i}&y_{fit,i}&z_{fit,i}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{3}$ can be calculated by:
@@
where $\theta_{i}=2{\pi}i/n,i=1,\cdots,n$. Then $z_{fit,i}$ can be obtained by combining $x_{fit,i},\ y_{fit,i}$ and the reconstructed point cloud of the target surface.
In order to avoid contact with the surface before milling starts, an offset should be set for all sampling points along their normal directions, where the normal vector of each point $\boldsymbol{n}_{i}\in\mathbb{R}^{3},i=1,\cdots,n$ can be calculated from the reconstructed point cloud. As a result, the 3D-fitted initial trajectory $\boldsymbol{p}_{ini}^{\{\text{Mi}\}}$ with respect to {Mi} can be obtained, where each element can be calculated by
@@
where $r_{d}$ is the radius of the drill tip, and $\Delta d$ is a manually set offset for safety concerns, taking into account both preliminary measurements of the 3D Surface Fitting module and a conservative margin to accommodate measurement uncertainty and possible unmodeled variations.
@@
The result of (7) is output to the Trajectory Planner (Section III-B) for the initialization of the milling trajectory.
III End-to-end Automation Workflow
With the integration of the proposed online pre-measurement pipeline (Section II), we improve the autonomous robotic bone micro-milling system and achieve an end-to-end automation workflow (see Fig. 2). In this section, we briefly introduce the intraoperative real-time feedback loop in the workflow, consisting of Completion Level Recognition (Section III-A) and Trajectory Planner (Section III-B), which are modified from the authors’ previous system [21] and ensure the system’s adaptive control after the online pre-measurement phase is completed and milling begins.
III-A Completion Level Recognition
The Completion Level Recognition aims to monitor the milling completion degree along the trajectory in real-time using a CNN based on the input RGB images. The CNN follows the same DSSD-inspired [3] architecture as the previous system [21], but is retrained on a new mouse skull dataset. The CNN outputs a bounding box for milling area detection and a heatmap image for pixel-wise completion level prediction. The network was trained on an augmented dataset derived from 587 manually collected and annotated images (a total of 18,784 training images) from videos of teleoperated robotic cranial window creation experiments conducted on 6 euthanized mice, achieving 77.6 mAP for the detection task and 24.32% MAPE for the completion prediction task, with a processing speed of 72 Hz. More training details (network architecture, dataset creation, data annotation, loss function, training condition, extended evaluation, and output example) are available in the Supplementary Materials.
It is important to note that although the 24.32 % MAPE may not seem ideal, this is due to the strictness of pixel-level completion prediction. In actual milling experiments, the system is relatively insensitive to variations at low completion levels and only needs to accurately capture the moment of penetration. The color and light response changes of the skull at penetration coincidentally enable the CNN to predict this critical moment with reasonable accuracy. Nevertheless, prediction errors can still occur, potentially causing experimental failures, which will be discussed in detail in Section IV-D3.
Through post-processing and sampling of the output image, an $n$-dimensional vector $\boldsymbol{c}$ can be obtained, where each element denotes the penetration percentage at the $i$-th sampling point along the real-time-updated trajectory, defined by
@@ III-A Completion Level Recognition
where i=1,2,,ni=1,2,...,n. Note that $c_{i}=0$ means untouched and $c_{i}=1$ means completely penetrated. $\boldsymbol{c}$ is output to the Trajectory Planner for real-time trajectory updates.
III-B Trajectory Planner
@@
The Trajectory Planner is responsible for initializing, updating and smoothing the milling trajectory. The updating process remains consistent with the previous system, employing a multimodal velocity damper to control the descent speed of the drill bit at the $i$-th sampling point to be proportionally reduced according to its completion level $c_{i}$, which is real-time output from the Completion Level Recognition and defined by (8).
@@
For the initialization process, in the previous system, due to the absence of direct surface measurements, the $z$-coordinates of all sampling points are initialized to 0 (see Fig. 5-(a)). In this study, with the integration of MSCS, a 3D-fitted initial trajectory $\boldsymbol{p}_{ini}$ is available, output from the 3D Surface Fitting (Section II-B) and defined by (1). The trajectory features a non-zero $z$-axis coordinate at each sampling point to pre-fit the uneven surface (see Fig. 5-(b)), enabling the microdrill to contact the target surface rapidly and spatially consistently.
@@
The smoothing process remains the same as in the previous system, where a constrained cubic spline interpolation method [8] is used to generate a continuous trajectory based on the $n$ discrete sampling points, thereby matching the high-precision encoders used in the robotic system [12]. Additionally, we hypothesize that if all $n$ sampling points are completely penetrated, the entire smoothed trajectory will be sufficiently milled. Therefore, when $c_{i}=1\ \ \forall i=1,\cdots,n$, a stop signal will be triggered and sent to the controller to automatically halt the robot’s movement and the drill’s rotation, marking the completion of the milling process. More details on the mathematical description of the Trajectory Planner are available in the Supplementary Materials.
IV Experiments
@@
In this section, experiments are first conducted to validate the effectiveness of the Automatic Calibration and the 3D Surface Fitting, which together constitute the proposed online pre-measurement pipeline. Subsequently, further experiments are conducted to evaluate the performance of the entire autonomous robotic bone micro-milling system. More details on the hardware setup and software implementation applied in the experiments are available in the Supplementary Materials.
IV-A Parameter selection
@@
To implement the proposed method for validations and experiments, the parameters mentioned in Section II and Section III are selected based on camera calibration, actual measurements, safety considerations, and ease of calculation: the number of trajectory sampling points $n=32$, the intrinsic parameters of cameras $c_{x}=480\ \mathrm{pixel}$, $c_{y}=270\ \mathrm{pixel}$, $P_{\rho x}=0.03947\ \mathrm{pixel/mm}$, $P_{\rho y}=0.03894\ \mathrm{pixel/mm}$, $d_{e}=0$, and $h_{\rho}=6.4\ \mathrm{pixel/mm}$, the ratio of the distance between the trajectory center and bregma to the distance between bregma and lambda $l=1/3$, the radius of the milling trajectory r=2mmr=2\ $\mathrm${mm}, the radius of the drill tip $r_{d}=0.5\ \mathrm{mm}$, the manually set offset $\Delta d=0.5\ \mathrm{mm}$, the number of calibration setpoints $k=5$ and their coordinates $\boldsymbol{p}_{1}^{\{\text{R}\}}=\begin{bmatrix}0&-20&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{2}^{\{\text{R}\}}=\begin{bmatrix}10&-20&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{3}^{\{\text{R}\}}=\begin{bmatrix}5&-10&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{4}^{\{\text{R}\}}=\begin{bmatrix}10&0&0\end{bmatrix}^{\text{T}},\ \boldsymbol{p}_{5}^{\{\text{R}\}}=\begin{bmatrix}0&0&0\end{bmatrix}^{\text{T}}\ \mathrm{mm}$. Additionally, the resolution of the images captured by the MSCS is set to $540\times 960$ to meet the requirement for each module.
IV-B Experiment 1: Validation of Automatic Calibration
@@
As described in Section II-A, the accuracy of Automatic Calibration is affected by the accuracy of 2D Keypoint Detection, 3D Reconstruction, and Landmark Calibration. Considering that the 3D Reconstruction depends on the MSCS, whose accuracy has been evaluated in prior research by the present authors [9], in Experiment 1, we focus on evaluating 2D Keypoint Detection and Landmark Calibration.
IV-B1 Training of 2D Keypoint Detection model
@@
A total of 3,294 images, each showing the drill tip within the field of view, were captured from videos of teleoperated cranial window experiments on 10 euthanized mice. No additional ethical approval was needed as the mice were reused from another study. For each image, a confidence map was generated by applying a 2D Gaussian to the manually annotated drill tip position, serving as ground truth (an example is shown in Fig. 6-(a)). Data augmentation, including blur, noise, color adjustment, geometric transformation, and resolution changes, expanded the dataset to 16,394 images, with 14,754 for training, 820 for validation, and 820 for testing.
@@
The models were trained for 30 epochs using the Adam optimizer with Mean Square Error (MSE) loss. The learning rate was initially set to $1\times 10^{-3}$, then reduced to $1\times 10^{-4}$ for epochs 11 to 20, and finally to $1\times 10^{-5}$ for epochs 21 to 30, with a batch size of 4. Training was implemented in PyTorch 1.8.0 with CUDA 11.1 on Python 3.8, running on Ubuntu 20.04 with an NVIDIA Quadro P6000 GPU.
IV-B2 Joint evaluation metrics
@@
Firstly, we randomly generate $m=20$ validation setpoints $\boldsymbol{p}_{val,i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,m$ within the reconstruction space of MSCS ($x\in[0,10],\ y\in[-20,0],\ z\in[0,10]\ \mathrm{mm}$) and send them to the robot so that the drill tip reaches these points. MSCS captures $m$ images each from left and right camera, outputting $m$ drill tip 2D positions each $\boldsymbol{p}_{l,i},\ \boldsymbol{p}_{r,i}\in\mathbb{R}^{2},i=1,\cdots,m$ via 2D Keypoint Detection (see Section II-A1). Then we manually annotate the left and right images, obtaining $m$ ground truths each $\boldsymbol{p}_{l,i}^{(gt)},\ \boldsymbol{p}_{r,i}^{(gt)}\in\mathbb{R}^{2},i=1,\cdots,m$. Through 3D Reconstruction (see Section II-A2), $m$ points of the 3D reconstructed drill tip with respect to {Mi} can be obtained. Then, by transforming these points back to {R} using the transformation calculated from Landmark Calibration (see Section II-A3), we can obtain $\boldsymbol{p}_{trans,i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,m$. The accuracies of 2D Keypoint Detection and Landmark Calibration are respectively evaluated by the Root Mean Square Error (RMSE) between $\boldsymbol{p}_{l,i},\ \boldsymbol{p}_{r,i}$ and $\boldsymbol{p}_{l,i}^{(gt)},\ \boldsymbol{p}_{r,i}^{(gt)}$, and between $\boldsymbol{p}_{trans,i}^{\{\text{R}\}}$ and $\boldsymbol{p}_{val,i}^{\{\text{R}\}}$.
Automatic Calibration mainly provides a basis for rapidly generating the initial surface-fitting trajectory, so its accuracy requirement prioritizes efficiency. Therefore, based on preliminary experiments and practical experience, the error in 2D keypoint detection is required not to exceed 10 pixels, and the error of Landmark Calibration is required not to exceed 1 mm.
IV-B3 Result and discussion
@@
The evaluation results, shown in Fig. 7, indicate an RMSE of 3.18 pixels for 2D Keypoint Detection and 0.77 mm for Landmark Calibration, both meeting the required accuracy. The validation setpoints and their reconstructed results are provided in the Supplementary Materials.
Specifically, as shown in Fig. 7-(a), all 20 validation points exhibited 2D detection errors within the threshold, demonstrating good accuracy and robustness of the model. For Landmark Calibration, although the 0.77 mm RMSE satisfies the requirement, it exceeds the safety offset $\Delta d$ (0.5 mm), which theoretically increases the risk of the drill contacting the skull surface before milling. Moreover, as shown in Fig. 7-(b), 4 points exceeded the error threshold. Further analysis revealed that these points were all located near the image periphery. This is likely due to the relatively large tilt angle ($\phi$), i.e., deviation from the vertical direction, which reduces the stability of 3D reconstruction in peripheral regions. In practice, trajectory sampling points are concentrated near the image center to ensure reconstruction accuracy. Among the 20 randomly generated validation points, 7 were located at the image periphery. After manually excluding these points, the RMSE of the remaining points decreased to 0.31 mm, which is below $\Delta d$, indicating that the actual calibration accuracy in the central region is significantly higher than the overall random-test performance. Future work will focus on improving 3D reconstruction accuracy by employing higher-resolution cameras and more robust models, thereby enhancing spatial localization performance.
IV-C Experiment 2: Validation of 3D Surface Fitting
As described in Section II-B, the accuracy of 3D Surface Fitting is affected by the accuracy of Trajectory Center Localization and 3D-fitted Initial Trajectory Generation. As the performance of 3D-fitted Initial Trajectory Generation is based on the reconstruction of the target surface, which has been evaluated in [9], and the transformation obtained via Landmark Calibration, which has been evaluated in Section IV-B, in Experiment 2, we focus on the evaluation of Trajectory Center Localization.
IV-C1 Training of CNN model
A dataset consisting of 523 images that satisfy the requirement that the mouse skulls in the images remain unmilled was captured from videos of the same 10 euthanized mice mentioned in Section IV-B, as well as 82 images from a publicly available dataset created by [23]. The creation of ground truths and data augmentation are the same as Section IV-B, resulting in a total of 8,774 images, with 7,019 for training, 877 for validation, and 878 for testing (an example is shown in Fig. 6-(b)). Training conditions and parameters are consistent to Section IV-B.
IV-C2 Metrics
Considering that the trajectory center is calculated by linearly weighting the detected bregma and lambda, with the detection errors of these two points considered independent, the error of Trajectory Center Localization can also be calculated by linearly weighting the detection RMSE of bregma and lambda. Same as Section IV-B, we require the error not to exceed 10 pixels.
IV-C3 Result and discussion